OpenStudy (anonymous):

how to draw r^2=4sin(theta) , in polar coordinates ?

OpenStudy (anonymous):

graphing calculator?

OpenStudy (anonymous):

Put in some simple \(\theta\) values like \(0,\pi/4,\pi/2\). Find corresponding \(r\) values. Then just connect the dots.

OpenStudy (anonymous):

Alternatively \[ r^3=4r\sin\theta\implies \left(\sqrt{x^2+y^2}\right)^3 = 4y \]If you think this is easier to graph, you can do that.

OpenStudy (anonymous):

but you are saying r^3=4r ... I didnt write an "r" on the right side

OpenStudy (jdoe0001):

solve for "r" and then grab a few random \(\theta\) 's to get the "r", plot away keep in mind that the "r" is the DEPENDENT variable

OpenStudy (anonymous):

i know how to choose values to find r .. but then what to do ?? i mean how to plot it on the polar coordinate system ??

OpenStudy (anonymous):

Use a ruler.

OpenStudy (anonymous):

figure out what scale you want to use.

OpenStudy (anonymous):

i dont know what i'm doing :( ....

OpenStudy (anonymous):

graphing polar curve actually isn't easy. You will need to treat it like it's in rectangular form, and use it to sort of guess how it would look like in polar coordinate this guy explained very well in his youtube videos http://www.youtube.com/watch?v=mDT_DG_A0JA

OpenStudy (anonymous):

thanxs that is a good explination , though I was just watching a similar video by the same guy Lol